At a "paint swap", Gayle found large supplies of Skylite Pink (12.5% red pigment)but only 3 gallons of MacIntosh Red (20% red). How many gallons of Skylite Pink should Gayle pick in order to mix Summer Rose (17% Red)?
Answer by Paul(988)

Let the amount of gallon needed be x:
Equation:
12.5x+3(20)=17(x+3)
12.5x+60=17x+51
4.5x=9
x=2
Hence, about 2 gallons is needed.
Paul.